# Are Large Language Models Memorizing Bug Benchmarks?

### v1

@@ VII Conclusion
In this paper, we measure data leakage risks across multiple widely used bug benchmarks and state-of-the-art open and closed-source models. Older benchmarks, especially Defects4J, exhibit higher memorization signals in models, particularly in smaller, older models with lower training budgets like codegen-multi 6B. Newer datasets such as Gitbug-Java, BugsInPy, BugsCpp show lower leakage risk, with similar NLL and 5-gram match to new 2024 repositories. We also find that newer models with higher training budgets display significantly lower risks of data leakage. Our findings suggest that researchers should consider both the models they use in evaluation and pair older benchmarks, such as Defects4J, with more recent benchmarks, such as Gitbug-Java, to minimize data contamination risk.
@@
(a) NLL ratios for LlaMa 3.1 8B.
(b) NLL ratios for Gemma 2B.
(c) NLL ratios for CodeLlaMa 7B.

### v2

@@ VII Conclusion
In this paper, we measure data leakage risks across multiple widely used bug benchmarks and state-of-the-art open-source models. Older benchmarks, especially Defects4J, exhibit higher memorization signals across all models, particularly in codegen-multi 6B. Newer datasets such as GitBug-Java, BugsInPy, BugsCpp show lower leakage risk, with similar NLL and 5-gram accuracy to new 2024 repositories. We also find that newer models with higher training budgets display significantly lower risks of data leakage. Our findings suggest that researchers should consider both the models they use in evaluation and pair older benchmarks, such as Defects4J, with more recent benchmarks, such as GitBug-Java, to improve model evaluation and address potential leakage issues appropriately.
@@
(c) NLL ratios for LlaMa 3.1 8B.
(d) NLL ratios for Gemma 2B.
(e) NLL ratios for StarCoder 7B.
(f) NLL ratios for CodeLLaMa 7B.
Figure 9: Heatmaps for NLL ratios
(a) Residual vs. Fitted Plot for average 5-gram accuracy.
(b) Residual vs. Fitted Plot for average NLL.
Figure 10: Residuals for the linear regressions.